OpenStudy (anonymous):

55% of what number is 70? Round to the nearest tenth.

hero (hero):

is/of = %/100 70/x = 55/100 55x = 7000 x = 7000/55

hero (hero):

x = 127.273

OpenStudy (anonymous):

127.27

OpenStudy (anonymous):

grrrr i was writing that :P

hero (hero):

lol

OpenStudy (anonymous):

how to write 82 1/2% of n as an expression